OpenStudy (anonymous):

I need to right a ratio for this situation in 3 ways. In greenville, About one in every eight homes has three computers.

OpenStudy (anonymous):

@Miracrown

Miracrown (miracrown):

There are a couple of different ways we can do this. Let's start by trying to express this information as a fraction.

Miracrown (miracrown):

Can you do that?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

Would the fraction possibly be 3 over 8

Miracrown (miracrown):

That is one of them.

Miracrown (miracrown):

Do you know how fractions relate to ratios?

OpenStudy (anonymous):

Ok so if its 3over8 then it would be 3:8 as another right?

Miracrown (miracrown):

Correct. :)

OpenStudy (anonymous):

Then just plainly 3 to 8 ?

OpenStudy (anonymous):

I guess my mind went blank for a moment.... xD

Miracrown (miracrown):

Yes, that's it! LOL don't worry it happens, at least you got the hang of it now. xD

OpenStudy (anonymous):

:)

Miracrown (miracrown):

(:

OpenStudy (anonymous):

i have another question on ratio's

OpenStudy (anonymous):

This time its taking 4 over 36 and telling me to write 3 ratios equal to it

Miracrown (miracrown):

Since ratios can be represented as fractions , we need 3 fractions that are the same as 4/36. Can you guess how we will find them?

OpenStudy (anonymous):

Maybe by dividing?

Miracrown (miracrown):

Close but not quite. Dividing will change the expression. I will give you a hint with a similar example.

Miracrown (miracrown):

\[\frac{ 1 }{ 2 } \space = \frac{ 3 }{ 9 } \space = \frac{ 11 }{33}\]

Miracrown (miracrown):

I have done this for 1/3 Does this give you any idea of how to handle 4/36?

OpenStudy (anonymous):

Not really sure, It looks like you took the 1 and the 3 and timed them by itself.

Miracrown (miracrown):

Sorry, that was a typo. It was supposed to be 1/3 and not 1/2. Let me fix it real quick

Miracrown (miracrown):

\[\frac{ 1 }{ 3 } \space = \frac{ 3 }{ 9 } \space = \frac{ 11 }{ 33 }\]

OpenStudy (anonymous):

Still a little confused here.

Miracrown (miracrown):

Ah, so good guess. The idea was to use simplified or un-simplified forms of the same fraction.

Miracrown (miracrown):

4/8 can be written as 2/4 , 8/16, and 1/2 as another example.

OpenStudy (anonymous):

alright that makes a little more sense, Let me look at see if i can figure it out.

Miracrown (miracrown):

Furthermore you can obtain these equivalent fractions by multiplying or dividing the numerator and the denominator by the same thing.

OpenStudy (anonymous):

1/8 2/16 and 3/24 are my answers at the moment

Miracrown (miracrown):

\[\frac{ 1 }{ 3 } = \space \frac{ 1 \space (3) }{ 3 \space (3) } \space = \frac{ 3 }{ 9 }\]

OpenStudy (anonymous):

That makes sense.

Miracrown (miracrown):

While those all are the same as 1/8 , 4/36 is not the same as those.

Miracrown (miracrown):

As a side note , if you had 4/32 instead of 4/36 your answer would be good.

OpenStudy (anonymous):

Alright.

OpenStudy (anonymous):

Would the top numbers stay the same or change to make my answer correct?

Miracrown (miracrown):

The numerators you mean? Ah, yes. It depends due to the infinitely many solutions , I can't give a clear answer for that. However , there do exist answers with those numerators.

OpenStudy (anonymous):

1/9 2/18 and 3/27??

Miracrown (miracrown):

Correct. (:

OpenStudy (anonymous):

Alright, Thank you.

Miracrown (miracrown):

yw

OpenStudy (anonymous):

That is all for now, I appreciate your help :)

Miracrown (miracrown):

You're most welcome. I am glad I could help =]